OpenStudy (anonymous):

Calc III problem: Find an equation of the plane that contains the y-axis and makes an angle of (pi/6) with the positive x-axis. I'm a little stuck. So far I understand that since the point contains the y-axis, then the plane should contain (0,0,0). After that, I'm sort of stuck. Thank you

ganeshie8 (ganeshie8):

start by finding the normal vector of plane

ganeshie8 (ganeshie8):

since the plane contains y axis, the y component of normal vector must be 0

OpenStudy (anonymous):

So the normal vector is (pi/6, 0, c)?

ganeshie8 (ganeshie8):

Also since the angle between normal vector and x axis is pi/6, the x component is same as the direction cosine : cos(pi/6) = sqrt(3)/2

OpenStudy (anonymous):

OHH the y component

ganeshie8 (ganeshie8):

the unit normal vector will be : \(\langle \frac{\sqrt{3}}{2}, 0, c \rangle\)

ganeshie8 (ganeshie8):

you can find the z component by setting the magnitude equal to 1

ganeshie8 (ganeshie8):

\(\large (\frac{\sqrt{3}}{2})^2 + 0^2 + c^2 = 1\) solve \(c\)

OpenStudy (anonymous):

And then you can use (0,0,0) as your point to solve for the equation of plane right?

ganeshie8 (ganeshie8):

Oh yes !

ganeshie8 (ganeshie8):

equation of plane with normal vector <a,b,c> is ax + by + cz = p

OpenStudy (anonymous):

Oh.. will it still work if I plug into the scalar equation form?

OpenStudy (anonymous):

like a(x-xnot) etc etc

ganeshie8 (ganeshie8):

it will work

OpenStudy (anonymous):

So should I get (sqrt(3)/2)x +0.5z = 0 ?

ganeshie8 (ganeshie8):

looks perfect ! u may multiply by 2 if want

OpenStudy (anonymous):

Thank you so much!

ganeshie8 (ganeshie8):

yw!